Page 367 Evaluate each of the following quadratic functions for the given
value of x. Show all your work.
f(x) =x^2 + 5x + 7; f(4)
g(x) =25 -x^2; g(-5)
h(x) = -x^2 +4x-1; h(-2)
f(x) =5x^2 -18;f(0)
g(x) =8x^2 -3x +10; g(2)
h(x) -3x^2 +15; h(10)

Answer by Alan3354(69443)

Evaluate each of the following quadratic functions for the given
value of x. Show all your work.
f(x) =x^2 + 5x + 7; f(4)
Sub 4 for x
f(4) = 4^2 + 5*4 + 7
f(4) = 16 + 20 + 7
f(4) = 43
--------------------
Do the same for the others.
g(x) =25 -x^2; g(-5)
h(x) = -x^2 +4x-1; h(-2)
f(x) =5x^2 -18;f(0)
g(x) =8x^2 -3x +10; g(2)
h(x) -3x^2 +15; h(10)